OpenStudy (anonymous):

which ordered pair is a solution of the equation y=-5x+4? (2,14) (5,-29) (3,-19) (5,-21)

OpenStudy (anonymous):

(2, 14) 14 = - 5(2) + 4 15 = -10 + 4 15 = -6 not a solution.. (5, - 29) - 29 = - 5(5) + 4 - 29 = - 25 + 4 - 29 = - 21 not a solution (3, - 19) - 19 = - 5(3) + 4 - 19 = - 15 + 4 - 19 = - 11 not a solution (5, - 21) - 21 = - 5(5) + 4 - 21 = - 25 + 4 - 21 = - 21 < ----- is the solution of the equation http://answers.yahoo.com/question/index?qid=20131010075552AAF1YyM